Where


C=300+0.2Y


I=550


G=800


X=400


M=100+0.4Y


T=0.1Y



Use the above information to answer the following questions



1) if C=300+0.2Yinterpret the meaning of the value 300


2)state the values of MPT,MPC and MPS on disposable income and MPM


3)calculate MPC and MPS on national income


4)find the level of national income in the economy


5) calculate net exports and interpret


6) if investment declines by 15% ,what will be the change in National income

1
Expert's answer
2022-01-29T11:06:11-0500


1)Y=C+G+I+Xn=300+800+550+400-100-0.4Y

1.4Y=1950

Y=1392.85



Y=C+G+I+Xn

C=300+0.2Y

Y=C+G+I+Xn=300+0.2Y+800+550+400=300+0.2(Y-0.1Y)+800+550+400-100-0.4Y=1950+0.2Y-0.02Y

Y-0.2Y+0.02Y+0.4Y=2050

1.22Y=1950

Y=1598.36


"C=300+0.2(Y-0.1Y)=300+0.2(1598.36-0.1\times1598.36)=587.70"


2) MPC=0.2

MPS=1-MPC=1-0.2=0.8

MRT=.0.1

MRM=0.4


3)MPC=0.2

MPS=1-MPC=1-0.2=0.8



4)"level of national income in the economy=Y-T=1598.36-1598.36\times0.1=1438.52"


5) "Xn=X-M=400-(100+0.4Y)=400-(100+0.4\times1598.36)=-339.34"



6)"Y=C+G+I+Xn=300+0.2(Y-0.1Y)+800+550-0.15\times550+400-100-0.4Y+400-100-0.4Y"

1.22Y=1482.5

Y=1215.16